# Why Foundation Models in Pathology Are Failing

### v2

@@ 1. Introduction
Foundation models (FMs) have transformed image and text processing by leveraging massive data corpora and self-supervised learning [1, 2, 3]. Their success in general computer vision inspired enthusiasm for analogous applications in histopathology, where the potential for automating diagnostic reasoning and multimodal retrieval is expected to be immense [4, 5]. However, emerging evidence indicates that pathology FMs underperform relative to expectations and exhibit critical weaknesses under realistic evaluation. This short paper analyzes these limitations and explores conceptual reasons behind their failures.
@@ 3. Why Are Pathology FMs Failing?
4. Synthesis: A Paradigm Mismatch
These failures reveal a mismatch between the foundation-model paradigm—built on abundance and homogeneity—and pathology, characterized by scarcity, heterogeneity, and context dependence. Scaling existing architectures without domain adaptation amplifies fragility rather than resolving it.
5. Ethical Considerations
The rapid proliferation of foundation model publications in pathology has created both excitement and concern within the scientific community. Alongside legitimate progress, several ethical challenges have emerged that undermine transparency, reproducibility, and trust in the field.
Journals and Editors – Editorial practices play a pivotal role in shaping scientific integrity. An emerging concern is the tendency of some journals or conference committees to invite non-experts or overly sympathetic reviewers, particularly in competitive or high-visibility areas such as foundation models. This can lead to insufficiently critical peer review and the publication of technically or conceptually weak work. A related issue is publication bias—the reluctance to publish “negative” or non-confirmatory results. Studies reporting low performance, instability, or bias in foundation models often face editorial resistance, skewing the literature toward overstated success and suppressing valuable evidence on limitations. This imbalance reinforces unrealistic expectations and slows methodological self-correction.
While empirical data show clear evidence of publication bias (i.e., fewer studies with negative or null results appear in the literature) [18], there is currently no robust published study demonstrating that journals routinely invite non-expert or sympathetic reviewers to favour a particular outcome. Nevertheless, the possibility remains plausible given structural incentives in AI/ML publishing and merits caution.
Researchers and Authors – At the research level, ethical responsibility extends beyond dataset curation and model evaluation to the transparency of reporting. Some studies may employ undisclosed training “tricks” or hyperparameter tuning strategies (omitting details), obscuring the true source of performance gains and limiting reproducibility. Similarly, non-rigorous or selectively chosen metrics—for example, reporting only accuracy and AUC values while ignoring F1 scores, calibration or fairness—can may present a misleading picture of model capability. Another recurring problem is potential “data leakage”, whether through overlapping patient samples, improper cross-validation, or hidden site bias. Such leakage can artificially inflate results, leading to spurious claims of generalization. In medical AI, where trust and safety are paramount, this constitutes not just methodological carelessness but an ethical lapse, as it risks translating flawed models into clinical claims.
Several meta-studies in machine-learning and medical-imaging have documented incomplete method disclosure (e.g., hyper-parameters, seeds, data-splits) and pervasive data-leakage biases (e.g., subject-wise leakage, tile reuse) that systematically inflate reported model performance [19, 20, 21, 22, 23].
Healthcare Management and Premature Adoption – A growing ethical risk lies beyond the research community: the premature adoption of foundation models by healthcare managers, administrators, or policy decision-makers eager to showcase “AI readiness” or institutional innovation. Driven by institutional pressure, marketing, or a desire to appear technologically progressive, some healthcare leaders may implement foundation models before their performance, generalizability, or bias have been adequately validated. Such enthusiasm, though often well-intentioned, can bypass the necessary clinical governance, risk assessment, and regulatory scrutiny. This overzealous adoption poses real dangers—misdiagnoses, inequitable treatment, and erosion of public trust—especially when AI systems are deployed under the illusion of universality or clinical maturity. Ethical stewardship therefore requires not only transparent science but also responsible decision-making at the administrative level, ensuring that AI implementation follows evidence, not hype.
Recent surveys of healthcare executives show strong optimism and investment in AI — for example, more than two-thirds of C-suite leaders report their organizations invest in AI, yet only  12-13% believe current algorithms are robust enough or have a clear strategy for clinical integration [24, 25, 26].
6. Conclusion
Foundation models have undeniably transformed vision and language, yet human tissue is not the Internet. The disappointing accuracy, fragility, and ethical blind spots of current pathology FMs reveal a deeper conceptual misalignment between data-driven generalization and the structured reasoning of medicine. Their challenges—ranging from architectural oversimplification and data scarcity to publication bias, hidden leakage, and premature adoption—underscore that progress in computational pathology cannot be achieved by scaling technology alone. True advancement will require rethinking the very foundations: developing models that see tissue as pathologists do—multi-scale, contextual, biologically grounded, and transparently validated within rigorous ethical and clinical frameworks. Only then can foundation models evolve from impressive demonstrations of computation to trustworthy instruments of medical understanding.

### v3

# Toward Clinically Grounded Foundation Models in Pathology

@@ 1. Introduction
Foundation models (FMs) have transformed image and text processing by leveraging massive data corpora and self-supervised learning [1, 2, 3]. Their success in general computer vision inspired enthusiasm for analogous applications in histopathology, where the potential for automating diagnostic reasoning and multimodal retrieval is expected to be immense [4, 5]. However, emerging evidence indicates that pathology FMs underperform relative to expectations and exhibit critical weaknesses under realistic evaluation. This short paper analyzes these limitations and investigates the conceptual factors contributing to their shortcomings, with the goal of informing future improvements in model design.
@@ 3. Understanding the FM Limitations
4. Conclusions
Foundation models have undeniably transformed vision and language, and their extension into medicine marks an important milestone in the evolution of artificial intelligence. Yet, the complexities of human tissue—and the interpretive reasoning that underpins pathology—demand more than scale alone. The limitations seen in current pathology FMs are signposts guiding the next phase of progress: from data-driven generalization toward models that integrate biological understanding, and clinical context.
Moving forward, genuine innovation will come from redefining the foundations themselves—designing systems that perceive tissue as pathologists do: multi-scale, contextual, and biologically grounded, validated within transparent and rigorous clinical frameworks. With such reorientation, foundation models can mature from computational feats into trusted instruments of discovery and diagnosis, bridging the gap between algorithmic power and medical insight.